# Investigating the Impact of Model Instability on Explanations and Uncertainty

### v1

@@ 1 Introduction
Due to this instability, it is difficult to know when we can trust a specific explanation. Ideally, we would like to use XAI to understand why a model succeeds and fails to identify points of failure in a model pipeline– these failures could arise from mistakes in the model training or ambiguity within the data. It is vital to understand when explanations are trustworthy, as the inclusion of XAI can cause an over-reliance on models Bauer et al. (2023); van der Waa et al. (2021), give users the false impression of global task understanding Chromik et al. (2021), and lead to overall poorer performance than if no human-AI collaboration Schmidt et al. (2020). Therefore, we would like to assess if the uncertainty of a model’s output can give any indication of an explanation’s quality, and if the instability of an explanation can provide insight into the model’s performance. We expect noise at inference time, especially for text data: Words can be accidentally ablated, mispelled or otherwise mutated. Different authors have distinct linguistic styles. New words emerge or change in meaning. Thanks to this noise, many SOTA language models suffer out-of-distribution issues and, thus, fail in real-world applications Alipanahi et al. (2022); Ribeiro et al. (2020). As large language models rely on drawing from large amounts of data (often stemming from sources with variable writing styles and formatting, like social media), we must understand how this “noise” in the data affects model’s performance, confidence, and explainability. As text perturbations can introduce some ambiguity into the data that is not present at training time, they should affect a model’s reported uncertainty alongside its explanation. Given the variety of language models available, it is also vital to compare how this relationship differs across different models and XAI methods.
In this paper, we conduct a large-scale empirical investigation into the effect of noise on Pre-trained Language Models, via a controlled experiment by artificially injecting varying degrees and types of realistic noise (see Table 1) and measuring the impact on model explanations and uncertainty. In this manner, we also investigate the relationship between explanation plausibility (the agreement between model saliency and ground-truth annotations) and model uncertainty. To assess if explanation instability reflects model instability, we limit our investigation to gradient-based techniques, rather than model approximation explanations. Here, we provide the following contributions:
  • We evaluate the relationship between uncertainty and explanation plausibility given perturbed and unperturbed data;
@@ 2 Related Work
Other works in the intersection of uncertainty and XAI quantify the uncertainty of a given explanation, by developing new models (Bykov et al., 2020) or looking at ensemble explanations Chai (2018); Slack et al. (2020); Marx et al. (2023), or they attempt to explain the causes of a model’s uncertainty Brown and Talbert (2022); Watson et al. (2023). In Marx et al. (2023), they find that the size of the dataset is inversely proportional to the uncertainty of the explanations, which suggests that, with increased training data, XAI techniques tend to converge and that epistemic uncertainty may affect XAI explanations. However, these methods do not look at existing links between XAI and uncertainty and look mainly at image and synthetic datasets.
Dataset Task Size SemEval 2013 Task 2 Sentiment Classification Training: 4133 Annotated Test: 1659 SST-2 + Hummingbird Sentiment Classification Training: 67349 Annotated Test: 62 HateXplain Hatespeech Detection Training: 15383 Annotated Test: 1142
@@
In summary, most studies investigating noise on model output look only at small levels of perturbation and focus on a small subset of language models (if any). Furthermore, they conflate different sources of uncertainty. In our paper, we investigate the effect of different scales of perturbations on a range of popular language models, including GPT2 and OPT. In addition, to avoid conflating sources of uncertainty, we use multiple measures of uncertainty to assess the relationship between model instability and saliency attribution.
@@ 3.4 Explanation techniques
We focus on local gradient-based explanations, which use backpropagation to compute a saliency heatmap over input features for a specific datapoint to audit a model’s decision. These explanations have been shown to perform best across many metrics, models, and tasks Atanasova et al. (2020), and, compared to perturbation-based techniques, like LIME, which approximate model performance, are model-centric and should give a more faithful representation of the instability within the model, rather than the technique. The simplest implementation uses the gradient of the input as the saliency score Simonyan et al. (2013); however, this can be very noisy Smilkov et al. (2017). Therefore, we rely on modified versions: SmoothGrad (SG) returns the average saliency map obtained by perturbing the original input with Gaussian noise Smilkov et al. (2017). Guided Backpropogation (GBP) uses a different computation of gradients (by ignoring all negative values) to visually improve its saliency maps Springenberg et al. (2014). InputXGradients (IXG) considers both the importance of the feature and the strength of the expressed dimension Shrikumar et al. (2016). IntegratedGradients (IG) accumulates the gradients between an input of interest and a neutral baseline Sundararajan et al. (2017). We use the Captum implementations Kokhlikyan et al. (2019).
@@ 4.1 Noise on uncertainty and explanations
4.1.1 The effect of perturbation hierarchy
@@
4.1.2 The effect of perturbation type
@@ 4.2 The relationship between uncertainty and explanation plausibility
Before Perturbation Including Perturbed Text Predictive uncertainty Epistemic uncertainty Predictive uncertainty Epistemic uncertainty dataset model GBP IXG IG SG GBP IXG IG SG GBP IXG IG SG GBP IXG IG SG SST-2 BERT 0.076 0.068 -0.128 -0.155 -0.052 -0.060 0.041 0.039 -0.104 -0.099 -0.069 -0.069 -0.240 -0.228 -0.248 -0.219 ELECTRA 0.040 0.002 -0.050 -0.089 -0.127 -0.065 -0.050 -0.058 -0.096 -0.096 -0.043 -0.050 -0.383 -0.380 -0.164 -0.175 RoBERTa 0.088 0.048 0.030 -$0$0.367 -0.330 -0.174 -0.200 -0.124 -0.101 -0.084 -0.069 -0.357 -0.324 -0.267 -0.246 GPT2 0.078 -0.033 0.124 -0.014 -0.150 -0.237 -0.036 -0.088 -0.092 -0.068 -0.013 -0.004 -0.232 -0.241 -0.094 -0.068 OPT -0.205 -0.205 -0.030 -0.030 -0.159 -0.159 -0.099 -0.099 -0.152 -0.152 -0.130 -0.130 -0.219 -0.219 -0.109 -0.109 SemEval BERT 0.237 0.248 0.238 0.249 0.235 0.247 0.234 0.247 0.149 0.165 0.150 0.165 0.151 0.166 0.148 0.164 ELECTRA 0.200 0.232 0.199 0.232 0.201 0.233 0.199 0.231 0.162 0.169 0.162 0.169 0.163 0.171 0.162 0.170 RoBERTa 0.213 0.234 0.212 0.234 0.215 0.235 0.213 0.235 0.149 0.155 0.148 0.154 0.149 0.155 0.147 0.153 GPT2 0.220 0.181 0.218 0.182 0.221 0.184 0.219 0.181 0.127 0.120 0.127 0.121 0.128 0.122 0.127 0.120 OPT 0.224 0.224 0.226 0.226 0.230 0.230 0.224 0.224 0.164 0.164 0.165 0.165 0.167 0.167 0.164 0.164 HateXplain BERT 0.268 0.270 0.265 0.262 0.211 0.229 0.263 0.267 0.293 0.178 0.297 0.181 0.243 0.139 0.259 0.148 ELECTRA 0.565 0.458 0.573 0.464 0.539 0.430 0.568 0.462 0.444 0.240 0.452 0.247 0.425 0.221 0.448 0.244 RoBERTa 0.529 0.434 0.517 0.424 0.502 0.407 0.503 0.408 0.396 0.218 0.390 0.213 0.371 0.195 0.379 0.201 GPT2 0.393 0.278 0.386 0.278 0.380 0.270 0.399 0.284 0.300 0.106 0.298 0.105 0.291 0.097 0.304 0.110 OPT 0.459 0.459 0.428 0.428 0.432 0.432 0.456 0.456 0.408 0.408 0.391 0.391 0.382 0.382 0.410 0.410
@@ 4.3 Robustness across perturbation type
In summary, while perturbation decreases model performance and explanation plausibility, it has a task and perturbation-dependent effect on uncertainty. Furthermore, high uncertainty of an output does not necessarily imply low explanation plausibility, as we find a moderate positive correlation between the measures on some datasets. Where uncertainty measures fail to align with model accuracy on some perturbation patterns, saliency map robustness can provide additional indication of model performance patterns; Integrated Gradients typically shows the greatest robustness to all types of noise; however, we can see model-specific patterns in susceptibility to adversarial perturbation.
@@ 5 Discussion
Across all models, realistic perturbations, such as butterfingers or synonym have the smallest impact on task performance and explanation plausibility, yet masking has the greatest impact. Furthermore, MASK has the greatest effect on both measures of uncertainty. Furthermore, when we look at model-level comparisons in Appendix C.2 and C.3, we find that the perturbations that have the maximal impact on model performance (e.g. MASK for ELECTRA in SemEval), also have a unique impact on saliency maps at low levels of perturbation, suggesting that the instability we see in explanations can provide some signal to model performance in the absence of labels. While SmoothGrad shows good all-around robustness to noise due to its regularization, it does not show specific patterns in model instability. In contrast, Integrated Gradients has relatively high robustness for smaller language models at low levels of perturbation and shows increased robustness to perturbations that minimally impact model performance (synonym and butterfinger). While lack of robustness is typically viewed as a deficiency of an XAI technique Hedström et al. (2023), we believe it can also be a signal of model instability: Epistemic uncertainty approximation measures work to perturb the model decision boundary to obtain a datapoint’s likelihood of class correspondence as an indicator of uncertainty. In contrast, perturbation, by introducing stochasticity in an input (much like SmoothGrad), also suggests the proximity of a datapoint to the decision boundary. In cases where popular uncertainty measures are prone to over-confidence, a lack of robustness may give some indication of uncertainty for a particular data point.
@@
In the case of hatespeech detection, UNK and l33t surprisingly reduce uncertainty (see Appendix C.2); this could explain the positive correlation between uncertainty and explanation plausibility for HateXplain, as highly perturbed examples will show lower uncertainty as plausibility decreases. The dataset is compiled from Twitter, and character substitutions may be used to hide potentially offensive terms. While there is no class difference regarding the proportion of words containing letters and numbers (0:0.695%, 1:0.975%, 2:0.912%), at manual inspection, we find examples of l33t-like speak in Classes 0 and 2 (e.g. h0e) that we do not find in the neutral class (e.g. WW2). The existence of these examples in the training data may have made the noise an indicator of a class, owing to the high “learnability” of this perturbation Zhang et al. (2022b). So, when noise is learned to be an indicator of class, uncertainty may show a positive correlation with output quality and explanation plausibility. However, we also see a weaker, positive relationship with the Twitter-based SemEval dataset, and we do not see saliency attributed to l33t noise in Appendix D.2; therefore, models trained with noisy data may show this positive relationship. This suggests that when these models have greater uncertainty, they may be more precise at identifying salient tokens. Other studies also suggest performance improvements after training models with noisy data Yu et al. (2024). We show in Appendix D.4 that, at very high perturbation, the strength of this relationship weakens (due to lack of meaningful tokens), but can remain weakly positive for simple tasks.
@@
Altogether, this suggests that the effect of perturbation on language models must be considered holistically across noise type and training data; realistic perturbations, which are expected to be more present in data drawn from social media, have a smaller impact on performance, uncertainty, and explanations. While uncertainty is not always a faithful indicator of local instability, weakened robustness to perturbations can provide additional information for model performance. Furthermore, as high uncertainty does not necessarily imply low explanation plausibility with noisy datasets, noise-augmented training may not only help model performance in out-of-domain tasks, but may also help ensure coherent explanations in low-confidence domains. For future work, we recommend the use of Integrated Gradients for smaller language models as it gives a more holistic depiction of model performance in adversarial conditions; however, as models scale in size, other gradient-based explanation techniques are more robust.
@@ A.2 SemEval, HateXplain
SemEval model BERT RoBERTa ELECTRA GPT2 OPT Learning rate 1e-5 1e-5 3e-6 8e-5 7e-6 Batch size 16 16 8 32 32 Epochs 3 3 5 7 1 Random seed 37 37 24 42 42 Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8 1e-8 Adam $\beta$ 1 0.9 0.9 0.9 0.9 0.9 Adam $\beta$ 2 0.999 0.999 0.999 0.999 0.999 LLRD None None None None None Decay type Linear Linear Linear Cosine Cosine Warmup Fraction 0 0 0 0.01 0.01 Attention Dropout 0.1 0.1 0.1 0.1 0.1 Dropout 0.1 0.1 0.1 0.1 0.1 Weight Decay 0 0 0 0.1 0.1 Test Accuracy 92% 94% 91% 91% 91%
@@
HateXplain model BERT RoBERTa ELECTRA GPT2 OPT Learning rate 2e-5 6e-6 2e-5 5e-5 9e-6 Batch size 32 32 8 32 8 Epochs 5 5 2 6 1 Random seed 2 2 6 42 42 Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8 1e-8 Adam $\beta$ 1 0.9 0.9 0.9 0.9 0.9 Adam $\beta$ 2 0.999 0.999 0.999 0.999 0.999 LLRD None None None None None Decay type Linear Linear Linear Cosine Cosine Warmup Fraction 0 0 0 0.01 0.01 Attention Dropout 0.1 0.1 0.1 0.1 0.1 Dropout 0.1 0.1 0.1 0.1 0.1 Weight Decay 0 0 0 0.1 0.1 Test Accuracy 68% 69% 70% 66% 69%

### v2

@@ 1 Introduction
Due to this instability, it is difficult to know when we can trust a specific explanation. Ideally, we would like to use XAI to understand both why a model succeeds and fails to identify points of failure in a model pipeline– these failures could arise from mistakes in the model training or ambiguity within the data. It is also vital to understand when explanations are trustworthy, as the inclusion of XAI can cause an over-reliance on models Bauer et al. (2023); van der Waa et al. (2021), give users the false impression of global task understanding Chromik et al. (2021), and lead to overall poorer performance than if no human-AI collaboration Schmidt et al. (2020). Therefore, we would like to assess if the uncertainty of a model’s output can give any indication of an explanation’s quality and if the instability of an explanation can provide insight into the model’s performance. We expect noise at inference time, especially for text: Words can be accidentally ablated, mispelled or otherwise mutated. Different authors have distinct linguistic styles, and new words emerge or change in meaning. Due to this noise, many SOTA language models suffer out-of-distribution issues and, thus, fail in real-world applications Alipanahi et al. (2022); Ribeiro et al. (2020). As large language models rely on drawing from large amounts of data (often stemming from sources with variable writing styles and formatting, like social media), we must understand how this “noise” in the data affects model’s performance, confidence, and explainability. As text perturbations can introduce some ambiguity into the data that is not present at training time, they should affect a model’s reported uncertainty alongside its explanation. Given the variety of language models available, it is also vital to compare how this relationship differs across different models and XAI methods.
In this paper, we conduct a large-scale empirical investigation into the effect of noise on Pre-trained Language Models via a controlled experiment by artificially injecting varying degrees and types of realistic noise (see Table 1) and measuring the impact on model explanations and uncertainty. In this manner, we also investigate the relationship between explanation plausibility (the agreement between model saliency and ground-truth annotations) and model uncertainty. To assess if explanation instability reflects model instability, we limit our investigation to gradient-based techniques, given their high-performance in robustness and plausibility measures Atanasova et al. (2020) and to limit additional uncertainty introduced by model approximation techniques, like LIME Zhang et al. (2019).
Here, we provide the following contributions:
  • We evaluate, for the first time, the relationship between uncertainty and explanation plausibility given perturbed and unperturbed data;
@@ 2 Related Work
Other works in the intersection of uncertainty and XAI quantify the uncertainty of a given explanation, by developing new models (Bykov et al., 2020) or looking at ensemble explanations Chai (2018); Slack et al. (2020); Marx et al. (2023), or they attempt to explain the causes of a model’s uncertainty Brown and Talbert (2022); Watson et al. (2023). In Marx et al. (2023), they find that the size of the dataset is inversely proportional to the uncertainty of the explanations, which suggests that, with increased training data, XAI techniques tend to converge; therefore, epistemic uncertainty may affect XAI explanations. However, these methods do not look at existing links between XAI and uncertainty and look mainly at image and synthetic datasets.
{tblr} hline1,5 = -0.08em, hline2 = -, Dataset & Task Size
SemEval 2013
Task 2 Sentiment
Classification Training: 4133
Annotated Test: 1659
SST-2 +
Hummingbird Sentiment
Classification Training: 67349
Annotated Test: 62
HateXplain Hatespeech
Detection Training: 15383
Annotated Test: 1142
@@
In summary, most studies investigating noise on model output look only at small levels of perturbation and focus on a small subset of language models. Furthermore, they conflate different sources of uncertainty in their investigation and do not assess their link to saliency attribution. In our paper, we investigate the effect of different scales of perturbations on a range of popular language models, including GPT2 and OPT. In addition, to avoid conflating sources of uncertainty, we use multiple measures of uncertainty to assess the relationship between model instability and saliency attribution.
@@ 3.4 Explanation techniques
We focus on local gradient-based explanations, which use backpropagation to compute a saliency heatmap over input features for a specific datapoint to audit a model’s decision. These explanations have been shown to perform best across many metrics, models, and tasks Atanasova et al. (2020), and, compared to perturbation-based techniques, like LIME, which approximate model performance, are model-centric and should give a more faithful representation of the instability within the model, rather than the technique Zhang et al. (2019). The simplest implementation uses the gradient of the input as the saliency score Simonyan et al. (2013); however, this can be very noisy Smilkov et al. (2017). Therefore, we rely on modified versions: SmoothGrad (SG) returns the average saliency map obtained by perturbing the original input with Gaussian noise Smilkov et al. (2017). Guided Backpropogation (GBP) uses a different computation of gradients (by ignoring all negative values) to visually improve its saliency maps Springenberg et al. (2014). InputXGradients (IXG) considers both the importance of the feature and the strength of the expressed dimension Shrikumar et al. (2016). IntegratedGradients (IG) accumulates the gradients between an input of interest and a neutral baseline Sundararajan et al. (2017). We use the Captum implementations Kokhlikyan et al. (2019).
@@ 4.1 Noise on uncertainty and explanations
The effect of perturbation hierarchy
@@
The effect of perturbation type
@@ 4.2 The relationship between uncertainty and explanation plausibility
{tblr} row3 = r, row4 = r, row5 = r, row6 = r, row7 = r, row8 = r, row9 = r, row10 = r, row11 = r, row12 = r, row13 = r, row14 = r, row15 = r, row16 = r, row17 = r, row18 = r, cell13 = c=8c, cell111 = c=8c, cell23 = c=4c, cell27 = c=4c, cell211 = c=4c, cell215 = c=4c, cell41 = r=5, cell91 = r=5, cell141 = r=5, vline4 = 1, vline4,8,12 = 2, vline3,7,11,15 = 3-18, hline1,19 = -0.08em, hline2-4 = -, hline9,14 = 2-18, columns = colsep=3pt & Before Perturbation Including Perturbed Text
Predictive uncertainty Epistemic uncertainty Predictive uncertainty Epistemic uncertainty
dataset model GBP IXG IG SG GBP IXG IG SG GBP IXG IG SG GBP IXG IG SG
SST-2 BERT 0.076 0.068 -0.128 -0.155 -0.052 -0.060 0.041 0.039 -0.104 -0.099 -0.069 -0.069 -0.240 -0.228 -0.248 -0.219
ELECTRA 0.040 0.002 -0.050 -0.089 -0.127 -0.065 -0.050 -0.058 -0.096 -0.096 -0.043 -0.050 -0.383 -0.380 -0.164 -0.175
RoBERTa 0.088 0.048 0.030 -$0$0.367 -0.330 -0.174 -0.200 -0.124 -0.101 -0.084 -0.069 -0.357 -0.324 -0.267 -0.246
GPT2 0.078 -0.033 0.124 -0.014 -0.150 -0.237 -0.036 -0.088 -0.092 -0.068 -0.013 -0.004 -0.232 -0.241 -0.094 -0.068
OPT -0.205 -0.205 -0.030 -0.030 -0.159 -0.159 -0.099 -0.099 -0.152 -0.152 -0.130 -0.130 -0.219 -0.219 -0.109 -0.109
SemEval BERT 0.237 0.248 0.238 0.249 0.235 0.247 0.234 0.247 0.149 0.165 0.150 0.165 0.151 0.166 0.148 0.164
ELECTRA 0.200 0.232 0.199 0.232 0.201 0.233 0.199 0.231 0.162 0.169 0.162 0.169 0.163 0.171 0.162 0.170
RoBERTa 0.213 0.234 0.212 0.234 0.215 0.235 0.213 0.235 0.149 0.155 0.148 0.154 0.149 0.155 0.147 0.153
GPT2 0.220 0.181 0.218 0.182 0.221 0.184 0.219 0.181 0.127 0.120 0.127 0.121 0.128 0.122 0.127 0.120
OPT 0.224 0.224 0.226 0.226 0.230 0.230 0.224 0.224 0.164  0.164 0.165 0.165 0.167 0.167 0.164 0.164
HateXplain BERT 0.268 0.270 0.265 0.262 0.211 0.229 0.263 0.267 0.293 0.178 0.297 0.181 0.243 0.139 0.259 0.148
ELECTRA 0.565 0.458 0.573 0.464 0.539 0.430 0.568 0.462 0.444 0.240 0.452 0.247 0.425 0.221 0.448 0.244
RoBERTa 0.529 0.434 0.517 0.424 0.502 0.407 0.503 0.408 0.396 0.218 0.390 0.213 0.371 0.195 0.379 0.201
GPT2 0.393 0.278 0.386 0.278 0.380 0.270 0.399 0.284 0.300 0.106 0.298 0.105 0.291 0.097 0.304 0.110
OPT 0.459 0.459 0.428 0.428 0.432 0.432 0.456 0.456 0.408 0.408 0.391 0.391 0.382 0.382 0.410 0.410
@@ 4.3 Robustness across perturbation type
In summary
While perturbation decreases model performance and explanation plausibility, it has a task and perturbation-dependent effect on uncertainty. Furthermore, high uncertainty of an output does not necessarily imply low explanation plausibility, as we find a moderate positive correlation between the measures on some datasets. Where uncertainty measures fail to align with model accuracy on some perturbation patterns, saliency map robustness can provide additional indication of model performance patterns; Integrated Gradients typically shows the greatest robustness to all types of noise; however, we can see model-specific patterns in susceptibility to adversarial perturbation.
@@ 5 Discussion
In this section, we discuss the causes behind patterns seen across the experiments in §4, which are further supplemented with extra analyses in our Appendix. We investigate over-arching patterns as well as patterns across perturbation types and datasets.
Overarching patterns
@@
Across all models, realistic perturbations, such as butterfingers or synonym have the smallest impact on task performance and explanation plausibility, yet masking has the greatest impact. Furthermore, MASK has the greatest effect on both measures of uncertainty. We expect that the embedding layer of the PLMs is better equipped to handle synonym-level perturbations, allowing the hidden representations of the input to change minimally, so that model performance and explanations are minimally impacted. Furthermore, we explore model-level differences in Appendix C.2 and C.3 and find that the perturbations that have the maximal impact on model performance (e.g. MASK for ELECTRA in SemEval) also uniquely impact saliency maps at low levels of perturbation. This suggests that the instability we see in explanations can provide some signal to model performance in the absence of labels. While SmoothGrad shows good all-around robustness to noise due to its regularization, it does not show specific patterns in model instability. In contrast, Integrated Gradients has relatively high robustness for smaller language models at low levels of perturbation and shows increased robustness to perturbations that minimally impact model performance (synonym and butterfinger). While lack of robustness is typically viewed as a deficiency of an XAI technique Hedström et al. (2023), we believe it can also be a signal of model instability: Epistemic uncertainty approximation measures work to perturb the model decision boundary to obtain a datapoint’s likelihood of class correspondence as an indicator of uncertainty. In contrast, perturbation, by introducing stochasticity in an input (much like SmoothGrad), also suggests the proximity of a datapoint to the decision boundary. In cases where popular uncertainty measures are prone to over-confidence, a lack of robustness may give some indication of uncertainty for a particular data point.
Dataset-level patterns
@@
The relationship between uncertainty and explanation plausibility after perturbation varies across datasets. For HateXplain, UNK and l33t surprisingly reduce uncertainty (see Appendix C.2); this could explain the positive correlation between uncertainty and explanation plausibility for the dataset, as highly perturbed examples will show lower plausibility, yet lower uncertainty. The dataset is compiled from Twitter, and character substitutions may hide potentially offensive terms. While we do not see a significant class difference regarding the proportion of words containing letters and numbers (0:0.695%, 1:0.975%, 2:0.912%), at manual inspection, we find examples of l33t-like speak in Classes 0 and 2 (e.g. ‘h0e’) that we do not find in the neutral class (e.g. ‘WW2’). The existence of these examples in the training data may have made the noise an indicator of a class, owing to the high “learnability” of this perturbation Zhang et al. (2022b), creating the positive correlation between uncertainty, output quality and explanation plausibility.
In Appendix D.2, we investigate if particular perturbation types are salient and find that saliency is not attributed to l33t noise, suggesting it is not a class indicator. Furthermore, we also see a weaker, positive relationship with the Twitter-based SemEval dataset. Therefore, though one would expect to see an inverse relationship between uncertainty and explanation plausibility, we only see this behaviour with the SST-2 dataset; we posit that models trained with noisy data instead show a positive relationship between uncertainty and explanation plausibility. When these models express greater uncertainty, they are more precise at identifying salient tokens, adding to other reported performance improvements after training models with noisy data Yu et al. (2024). We show in Appendix D.4 that, at very high perturbation, the strength of this relationship weakens (due to lack of meaningful tokens), but can remain weakly positive for simple tasks.
In summary
@@
The results suggest that the effect of perturbation on language models must be considered holistically across noise type and training data; realistic perturbations, like synonyms and mispellings, which are expected to be more prevalent in social media, have a smaller impact on performance, uncertainty, and explanations. While uncertainty is not always a faithful indicator of local instability, weakened robustness to perturbations can provide additional information for model performance. Furthermore, as high uncertainty does not necessarily imply low explanation plausibility with noisy datasets, noise-augmented training may not only help model performance in out-of-domain tasks, but may also help ensure coherent explanations in low-confidence domains. Further research is required to devise a metric to estimate explanation quality at a datapoint-level, as current uncertainty measures are not reflective of explanation quality. For future work, we recommend the use of Integrated Gradients for smaller language models as it gives a more holistic depiction of model performance in adversarial conditions; however, as models scale in size, other gradient-based explanation techniques are more robust.
@@ A.2 SemEval, HateXplain
{tblr} row1 = c, cell11 = c=6, vline2 = 2-16, hline1,17 = -0.08em, hline2 = -0.05em, hline3 = 1-50.03em, hline3 = 6, columns = colsep=2pt, SemEval &
model BERT RoBERTa ELECTRA GPT2 OPT
Learning rate 1e-5 1e-5 3e-6 8e-5 7e-6
Batch size 16 16 8 32 32
Epochs 3 3 5 7 1
Random seed 37 37 24 42 42
Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8 1e-8
Adam $\beta$ 1 0.9 0.9 0.9 0.9 0.9
Adam $\beta$ 2 0.999 0.999 0.999 0.999 0.999
LLRD None None None None None
Decay type Linear Linear Linear Cosine Cosine
Warmup
Fraction 0 0 0 0.01 0.01
Attention
Dropout 0.1 0.1 0.1 0.1 0.1
Dropout 0.1 0.1 0.1 0.1 0.1
Weight Decay 0 0 0 0.1 0.1
Test Accuracy 92% 94% 91% 91% 91%
@@
{tblr} row1 = c, cell11 = c=6, vline2 = 2-16, hline1,17 = -0.08em, hline2 = -0.05em, hline3 = 1-50.03em, hline3 = 6, columns = colsep=2pt, HateXplain &
model BERT RoBERTa ELECTRA GPT2 OPT
Learning rate 2e-5 6e-6 2e-5 5e-5 9e-6
Batch size 32 32 8 32 8
Epochs 5 5 2 6 1
Random seed 2 2 6 42 42
Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8 1e-8
Adam $\beta$ 1 0.9 0.9 0.9 0.9 0.9
Adam $\beta$ 2 0.999 0.999 0.999 0.999 0.999
LLRD None None None None None
Decay type Linear Linear Linear Cosine Cosine
Warmup
Fraction 0 0 0 0.01 0.01
Attention
Dropout 0.1 0.1 0.1 0.1 0.1
Dropout 0.1 0.1 0.1 0.1 0.1
Weight Decay 0 0 0 0.1 0.1
Test Accuracy 68% 69% 70% 66% 69%